OpenStudy (anonymous):

HOW DO I SOLVE THE FOLLOWING: 1/3X-1/6X=2/5-7/10 AND -2(X-1)-3(2X+1)=-9

OpenStudy (anonymous):

Are those two independent equations?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

\[\left(\frac{1}{3}\right)X-\left(\frac{1}{6}\right)X=\frac{2}{5}-\frac{7}{10} \]Combine the terms on each side of the equation.\[\frac{X}{6}=-\frac{3}{10} \]Multiply each side by 30, a number which is divisible by 6 and 10 and then simplify the result.\[30\frac{X}{6}=-\frac{3}{10}30 \]\[5 X=-9 \]\[X=\frac{-9}{5} \]Second equation.\[-2(X-1)-3(2X+1)=-9 \]Perform the multiplications in the LHS.\[-2X+2-6X-3=-9\]Combine similar terms on the LHS.\[-1-8 X=-9\]\[X=1 \]